OpenStudy (readergirl12):

Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number? I don't understand..

OpenStudy (anonymous):

3x−7=9+5x solve for x x = ??

OpenStudy (readergirl12):

Could you tell me how?

OpenStudy (readergirl12):

Oh..Thank you. c: